OpenStudy (anonymous):

y = 55x. Use the x values for the number of hours: 0, 10, and 20

OpenStudy (anonymous):

here you have to first plug in 0 for x in the equation.Evaluate. Then, plug in 10 for x in the equation. Evluate Finally plug in 20 for x and evluate

OpenStudy (anonymous):

you will get 3 seperate answers

OpenStudy (anonymous):

so will it be {(0, 0), (10, 55), (1100, 20)} ?

OpenStudy (anonymous):

not quite: (0,0), (10,550),(20,1210)

OpenStudy (anonymous):

whoop!

OpenStudy (anonymous):

yeah, y=55(0)=0 y=55(10)=550 y=55(20)=1210